# PSDNorm: Test-Time Temporal Normalization for Deep Learning in Sleep Staging

### v2

@@ Acknowledgements
This work was supported by the grants ANR-22-PESN-0012 to AC under the France 2030 program, ANR-20-CHIA-0016 and ANR-20-IADJ-0002 to AG while at Inria, and ANR-23-ERCC-0006 to RF, all from Agence nationale de la recherche (ANR). This project has also received funding from the European Union’s Horizon Europe research and innovation programme under grant agreement 101120237 (ELIAS).
@@
Numerical computation was enabled by the scientific Python ecosystem: Matplotlib [50], Scikit-learn [51], Numpy [52], Scipy [53], PyTorch [48] and MNE [44].
@@ A.1 Proof of the Bures-Wasserstein geodesic (6) between covariance matrices of structure (2)
From [32], the geodesic associated with the Bures-Wasserstein metric between two covariance matrices ${\boldsymbol{\Sigma}}^{(s)}$ and ${\boldsymbol{\Sigma}}^{(t)}$ is given by
@@ A.2 Balanced datasets
In the main paper, we report results across different training set sizes. Since the datasets are highly imbalanced (e.g., ABC has 44 subjects, SHHS has 5,730), we create balanced subsets by randomly selecting up to $N$ subjects per dataset. This avoids over-representing the largest dataset and ensures greater diversity in the training data. We consider four values of $N$: 40, 100, 200, and 400. The average number of subjects in each balanced set is shown in Table 3. Notably, the balanced set with 400 subjects contains roughly four times less data than the full dataset.
@@ A.3 U-Time: CNN for time series segmentation
U-Time [45, 15] is a convolutional neural network (CNN) inspired by the U-Net architecture [54], designed for segmenting temporal sequences. U-Time maps sequential inputs of arbitrary length to sequences of class labels on a freely chosen temporal scale. The architecture is composed of several encoder and decoder blocks, with skip connections between them.
@@
The decoding part of U-Time is symmetrical to the encoding part. Each decoder block doubles the signal length and decreases the number of channels. It is composed of a convolutional layer, an activation function, a BatchNorm layer, an upsampling layer and a concatenation layer of the skip connection of the corresponding encoding block. Finally, the segmentation head applies two convolutional layers with an activation function in between to output the final segmentation. It should be noted that U-Time employs BatchNorm layers but other normalization layers, such as LayerNorm [7] or InstanceNorm [8] are possible.
@@
The architecture is inspired from Braindecode [55]. The implementation is improved to make it more efficient and faster. One epoch of training takes about 30 min on a single H100 GPU.
Architecture: CNNTransformer
The CNNTransformer is a hybrid architecture designed for multichannel time series classification inspired by transformers for EEG-Data [20, 21, 46, 23]. It combines convolutional feature extraction with long-range temporal modeling via a Transformer encoder at epoch-level. The model processes an input tensor of shape $(B,S,C,T)$, where $B$ is the batch size, $S$ is the number of temporal segments, $C$ is the number of input channels, and $T$ is the number of time samples per segment. It outputs a tensor of shape $(B,n_{\text{classes}},S)$, where $n_{\text{classes}}$ is the number of classes and $S$ is the number of epochs.
@@
Transformer Encoder Multi-head attention $2$ Transformer layers with $d_{\text{model}}$ embedding dimension, $n_{\text{head}}$ heads, and feedforward sublayers $(B,d_{\text{model}},S)$
@@
A.4 Equation for BatchNorm and InstanceNorm
@@
The BatchNorm layer [6] normalizes features maps in a neural network to have zero mean and unit variance. At train time, given a batch =$\mathcal{B}=\{\mathbf{G}^{(1)},\dots,\mathbf{G}^{(N)}\}\subset\mathbb{R}^{c\times\ell}$ of $N$ pre-BatchNorm feature maps and for all $j,m,l\in\left\llbracket 1,N\right\rrbracket\,\times\left\llbracket 1,c\right\rrbracket\,\times\left\llbracket 1,\ell\right\rrbracket\,$, the BatchNorm layer is computed as
@@
Another popular normalization is the InstanceNorm layer [8]. During training, InstanceNorm operates similarly to (14), but the mean and variance are computed per sample instead of across the batch dimension, i.e., $\widehat{\mu}_{m}^{(j)}$ and $\widehat{\sigma}_{m}^{(j)}$ are computed for each sample $j$,
@@
A.5 F1 score vs. Balanced Accuracy
In the main paper, we report Balanced Accuracy scores, which account for class imbalance in sleep stage classification. Prior work, such as the U-Time paper [45], uses the F1 score to evaluate performance. In Table 5, we report F1 scores on the left-out datasets. These scores are slightly higher than the Balanced Accuracy scores and are comparable to those reported in the U-Time paper.
@@
BatchNorm LayerNorm InstanceNorm PSDNorm(F=5) PSDNorm(F=9) PSDNorm(F=17)
All subjects ABC 79.80 $\pm$ 0.34 77.86 $\pm$ 0.80 78.36 $\pm$ 1.20 78.08 $\pm$ 0.78 76.93 $\pm$ 1.29 77.44 $\pm$ 0.58
CCSHS 88.32 $\pm$ 0.49 87.22 $\pm$ 0.51 88.73 $\pm$ 0.52 88.79 $\pm$ 0.99 87.86 $\pm$ 1.08 88.91 $\pm$ 0.28
CFS 87.01 $\pm$ 0.18 85.61 $\pm$ 0.16 87.62 $\pm$ 0.27 87.06 $\pm$ 0.77 87.26 $\pm$ 0.42 86.52 $\pm$ 0.99
CHAT 66.56 $\pm$ 1.42 61.32 $\pm$ 2.25 64.19 $\pm$ 4.63 71.86 $\pm$ 0.95 68.61 $\pm$ 5.36 71.12 $\pm$ 1.09
HOMEPAP 76.20 $\pm$ 1.25 76.15 $\pm$ 1.13 77.66 $\pm$ 0.58 77.85 $\pm$ 1.29 77.57 $\pm$ 0.68 77.18 $\pm$ 1.25
MASS 76.06 $\pm$ 1.69 73.95 $\pm$ 5.80 76.94 $\pm$ 1.12 77.16 $\pm$ 1.73 78.50 $\pm$ 1.82 76.86 $\pm$ 4.06
MROS 83.69 $\pm$ 0.39 82.22 $\pm$ 1.27 83.95 $\pm$ 0.53 83.51 $\pm$ 0.84 84.63 $\pm$ 1.02 83.96 $\pm$ 0.87
PhysioNet 76.26 $\pm$ 1.27 70.40 $\pm$ 0.14 73.84 $\pm$ 0.93 73.51 $\pm$ 3.05 73.80 $\pm$ 1.24 74.76 $\pm$ 0.50
SHHS 76.98 $\pm$ 0.70 75.98 $\pm$ 0.22 79.12 $\pm$ 0.96 79.26 $\pm$ 1.35 78.65 $\pm$ 1.04 77.69 $\pm$ 0.88
SOF 85.49 $\pm$ 0.58 84.23 $\pm$ 1.30 85.50 $\pm$ 0.86 84.14 $\pm$ 1.05 85.54 $\pm$ 0.16 85.33 $\pm$ 0.16
Mean 79.64 $\pm$ 0.41 77.57 $\pm$ 0.73 79.59 $\pm$ 0.25 80.12 $\pm$ 0.57 79.94 $\pm$ 0.80 79.98 $\pm$ 0.47
Balanced@400 ABC 81.00 $\pm$ 0.11 79.50 $\pm$ 0.49 80.56 $\pm$ 0.39 81.12 $\pm$ 0.37 80.80 $\pm$ 0.22 80.90 $\pm$ 0.31
CCSHS 89.83 $\pm$ 0.19 89.01 $\pm$ 0.43 89.39 $\pm$ 0.16 89.13 $\pm$ 0.17 89.22 $\pm$ 0.17 89.45 $\pm$ 0.57
CFS 88.30 $\pm$ 0.52 87.39 $\pm$ 0.06 88.45 $\pm$ 0.17 88.52 $\pm$ 0.15 88.32 $\pm$ 0.34 88.23 $\pm$ 0.47
CHAT 65.77 $\pm$ 4.06 65.25 $\pm$ 3.96 71.35 $\pm$ 2.75 72.16 $\pm$ 2.21 71.93 $\pm$ 3.01 72.99 $\pm$ 4.01
HOMEPAP 77.06 $\pm$ 0.14 76.62 $\pm$ 1.06 77.50 $\pm$ 0.46 77.30 $\pm$ 0.24 77.78 $\pm$ 0.62 77.98 $\pm$ 0.90
MASS 77.27 $\pm$ 1.42 74.21 $\pm$ 2.05 75.12 $\pm$ 2.08 76.00 $\pm$ 3.00 74.16 $\pm$ 4.21 74.10 $\pm$ 0.47
MROS 85.53 $\pm$ 0.48 84.02 $\pm$ 0.95 85.22 $\pm$ 0.19 85.02 $\pm$ 0.42 85.37 $\pm$ 0.30 84.93 $\pm$ 0.53
PhysioNet 74.98 $\pm$ 1.84 74.29 $\pm$ 1.50 75.07 $\pm$ 1.05 75.29 $\pm$ 1.21 75.41 $\pm$ 0.68 74.57 $\pm$ 1.16
SHHS 78.95 $\pm$ 0.92 78.04 $\pm$ 1.21 80.30 $\pm$ 1.29 80.32 $\pm$ 0.91 79.22 $\pm$ 1.27 80.06 $\pm$ 1.30
SOF 86.30 $\pm$ 0.40 85.82 $\pm$ 0.22 86.57 $\pm$ 0.60 86.99 $\pm$ 0.33 86.75 $\pm$ 0.54 86.76 $\pm$ 0.23
Mean 80.50 $\pm$ 0.51 79.41 $\pm$ 0.73 80.95 $\pm$ 0.36 81.19 $\pm$ 0.11 80.89 $\pm$ 0.00 81.00 $\pm$ 0.28
A.6 Impact of Whitening and Target Covariance
@@
BatchNorm InstanceNorm PSDNorm(F=5) PSDNorm(F=9) PSDNorm(F=17)
Whitening Barycenter Whitening Barycenter Whitening Barycenter Whitening
ABC 78.26 $\pm$ 1.33 78.73 $\pm$ 0.42 78.18 $\pm$ 0.68 77.86 $\pm$ 1.33 78.18 $\pm$ 0.91 78.16 $\pm$ 0.93 77.76 $\pm$ 1.00 78.34 $\pm$ 0.33
CCSHS 87.42 $\pm$ 0.16 87.62 $\pm$ 0.42 87.58 $\pm$ 0.30 87.80 $\pm$ 0.23 87.35 $\pm$ 0.52 87.53 $\pm$ 0.21 87.62 $\pm$ 0.48 87.90 $\pm$ 0.57
CFS 84.32 $\pm$ 0.57 84.72 $\pm$ 0.33 84.29 $\pm$ 0.36 84.01 $\pm$ 0.60 84.06 $\pm$ 0.10 84.20 $\pm$ 0.23 84.46 $\pm$ 0.06 84.04 $\pm$ 0.91
CHAT 66.55 $\pm$ 0.88 64.43 $\pm$ 4.41 70.28 $\pm$ 1.70 69.07 $\pm$ 3.73 68.11 $\pm$ 3.94 69.31 $\pm$ 2.50 69.88 $\pm$ 0.46 69.43 $\pm$ 1.80
HOMEPAP 75.25 $\pm$ 0.50 76.47 $\pm$ 0.63 76.83 $\pm$ 0.61 76.13 $\pm$ 0.93 76.61 $\pm$ 0.74 76.45 $\pm$ 0.29 76.49 $\pm$ 0.45 76.81 $\pm$ 0.22
MASS 70.00 $\pm$ 1.91 71.52 $\pm$ 1.13 72.77 $\pm$ 1.09 69.11 $\pm$ 1.51 73.07 $\pm$ 1.30 69.01 $\pm$ 0.78 72.23 $\pm$ 2.40 70.41 $\pm$ 1.50
MROS 80.37 $\pm$ 0.20 80.28 $\pm$ 0.21 80.26 $\pm$ 0.11 80.50 $\pm$ 0.75 80.32 $\pm$ 0.22 80.52 $\pm$ 0.15 80.70 $\pm$ 0.42 79.96 $\pm$ 0.53
PhysioNet 75.81 $\pm$ 0.13 74.68 $\pm$ 0.55 74.82 $\pm$ 2.11 74.58 $\pm$ 1.57 73.77 $\pm$ 1.73 75.08 $\pm$ 2.18 75.09 $\pm$ 0.97 75.61 $\pm$ 0.71
SHHS 76.44 $\pm$ 0.92 78.68 $\pm$ 0.37 78.88 $\pm$ 0.68 78.77 $\pm$ 0.67 77.28 $\pm$ 0.91 78.85 $\pm$ 0.16 78.41 $\pm$ 0.49 $7$$\pm$ 1.35
SOF 81.08 $\pm$ 1.14 80.68 $\pm$ 1.38 79.49 $\pm$ 0.41 80.10 $\pm$ 0.62 81.44 $\pm$ 0.97 80.53 $\pm$ 0.05 81.07 $\pm$ 0.66 79.63 $\pm$ 0.89
Mean 77.55 $\pm$ 0.34 77.78 $\pm$ 0.46 78.34 $\pm$ 0.42 77.79 $\pm$ 0.30 78.02 $\pm$ 0.67 77.96 $\pm$ 0.12 78.37 $\pm$ 0.38 77.99 $\pm$ 0.35
A.7 Generalization of PSDNorm in CNNTransformer

### v3

@@ Acknowledgements
This work was supported by the grants ANR-22-PESN-0012 to AC under the France 2030 program, ANR-20-CHIA-0016 and ANR-20-IADJ-0002 to AG while at Inria, and ANR-23-ERCC-0006, ANR-25-PEIA-0005 and ANR-23-IACL-0005 to RF, all from Agence nationale de la recherche (ANR). This work is supported by Hi! PARIS and ANR/France 2030 program (ANR-23-IACL-0005). This project has also received funding from the European Union’s Horizon Europe research and innovation programme under grant agreement 101120237 (ELIAS).
@@
Numerical computation was enabled by the scientific Python ecosystem: Matplotlib Hunter (2007), Scikit-learn Pedregosa et al. (2011), Numpy Harris et al. (2020), Scipy Virtanen et al. (2020), PyTorch Paszke et al. (2019) and MNE Gramfort et al. (2013).
@@ A.1 Proof of the Bures-Wasserstein geodesic (6) between covariance matrices of structure (2)
From Bhatia et al. (2019), the geodesic associated with the Bures-Wasserstein metric between two covariance matrices ${\boldsymbol{\Sigma}}^{(s)}$ and ${\boldsymbol{\Sigma}}^{(t)}$ is given by
@@ A.2 Balanced datasets
In the main paper, we report results across different training set sizes. Since the datasets are highly imbalanced (e.g., ABC has 44 subjects, SHHS has 5,730), we create balanced subsets by randomly selecting up to $N$ subjects per dataset. This avoids over-representing the largest dataset and ensures greater diversity in the training data. We consider four values of $N$: 40, 100, 200, and 400. The average number of subjects in each balanced set is shown in Table˜3. Notably, the balanced set with 400 subjects contains roughly four times less data than the full dataset.
@@ A.3 U-Time: CNN for time series segmentation
U-Time Perslev et al. (2019; 2021) is a convolutional neural network (CNN) inspired by the U-Net architecture Ronneberger et al. (2015), designed for segmenting temporal sequences. U-Time maps sequential inputs of arbitrary length to sequences of class labels on a freely chosen temporal scale. The architecture is composed of several encoder and decoder blocks, with skip connections between them.
@@
The decoding part of U-Time is symmetrical to the encoding part. Each decoder block doubles the signal length and decreases the number of channels. It is composed of a convolutional layer, an activation function, a BatchNorm layer, an upsampling layer and a concatenation layer of the skip connection of the corresponding encoding block. Finally, the segmentation head applies two convolutional layers with an activation function in between to output the final segmentation. It should be noted that U-Time employs BatchNorm layers but other normalization layers, such as LayerNorm Ba et al. (2016) or InstanceNorm Ulyanov (2016) are possible.
@@
The architecture is inspired from Braindecode Schirrmeister et al. (2017). The implementation is improved to make it more efficient and faster. One epoch of training takes about 30 min on a single H100 GPU.
A.4 Architecture: CNNTransformer
The CNNTransformer is a hybrid architecture designed for multichannel time series classification inspired by transformers for EEG-Data Wang et al. (2024); Phan et al. (2022b); Yang et al. (2023); Thapa et al. (2025). It combines convolutional feature extraction with long-range temporal modeling via a Transformer encoder at epoch-level. The model processes an input tensor of shape $(B,S,C,T)$, where $B$ is the batch size, $S$ is the number of temporal segments, $C$ is the number of input channels, and $T$ is the number of time samples per segment. It outputs a tensor of shape $(B,n_{\text{classes}},S)$, where $n_{\text{classes}}$ is the number of classes and $S$ is the number of epochs.
@@ A.4 Architecture: CNNTransformer
Transformer Encoder Multi-head attention 22 Transformer layers with $d_{\text{model}}$ embedding dimension, $n_{\text{head}}$ heads, and feedforward sublayers $(B,d_{\text{model}},S)$
@@
A.5 Equation for BatchNorm and InstanceNorm
@@
The BatchNorm layer Ioffe and Szegedy (2015) normalizes features maps in a neural network to have zero mean and unit variance. At train time, given a batch $\mathcal{B}=\{\mathbf{G}^{(1)},\dots,\mathbf{G}^{(N)}\}\subset\mathbb{R}^{c\times\ell}$ of $N$ pre-BatchNorm feature maps and for all $j,m,l\in\left\llbracket 1,N\right\rrbracket\,\times\left\llbracket 1,c\right\rrbracket\,\times\left\llbracket 1,\ell\right\rrbracket\,$, the BatchNorm layer is computed as
@@
Another popular normalization is the InstanceNorm layer Ulyanov (2016). During training, InstanceNorm operates similarly to (14), but the mean and variance are computed per sample instead of across the batch dimension, i.e., $\widehat{\mu}_{m}^{(j)}$ and $\widehat{\sigma}_{m}^{(j)}$ are computed for each sample $j$,
@@
A.6 Sensitivity to Filter Size
Figure 6: Performance of PSDNorm with varying filter sizes. The BACC score is plotted against the filter size used with U-Sleep.
The filter size $f$ in PSDNorm controls the temporal context used for normalization, influencing the strength of adaptation to temporal variations. Figure˜6 shows the impact of different $f$ values on the BACC score across datasets using U-Sleep trained on balanced@400. This experiments shows that for any $f$, PSDNorm consistently improves performance compared to other normalization techniques. Taking a $f$ between 5 and 11 yields the best results, with a peak at $f=5$. Smaller values (e.g., $f=1$, equivalent to InstanceNorm) provide less adaptation, while larger values (e.g., $f=21$) may over-smooth temporal variations, leading to diminished performance. Overall, the results shows that $f$ is not so sensitive yielding good performance for a wide range of values.
A.7 F1 score vs. Balanced Accuracy
In the main paper, we report Balanced Accuracy scores, which account for class imbalance in sleep stage classification. Prior work, such as the U-Time paper Perslev et al. (2019), uses the F1 score to evaluate performance. In Table 5, we report F1 scores on the left-out datasets. These scores are slightly higher than the Balanced Accuracy scores and are comparable to those reported in the U-Time paper.
@@ A.6 Sensitivity to Filter Size
Dataset BatchNorm LayerNorm InstanceNorm TMA PSDNorm(F=5)
All subjects ABC $81.00_{\pm 0.11}$ $79.50_{\pm 0.49}$ $80.56_{\pm 0.39}$ $80.89_{\pm 0.06}$ $\mathbf{81.12_{\pm 0.37}}$
CCSHS $\mathbf{89.83_{\pm 0.19}}$ $89.01_{\pm 0.43}$ $89.39_{\pm 0.16}$ $89.37_{\pm 0.11}$ $89.13_{\pm 0.17}$
[…]
CHAT $65.77_{\pm 4.06}$ $65.25_{\pm 3.96}$ $71.35_{\pm 2.75}$ $71.80_{\pm 2.66}$ $\mathbf{72.16_{\pm 2.21}}$
[…]
MROS $\mathbf{85.53_{\pm 0.48}}$ $84.02_{\pm 0.95}$ $85.22_{\pm 0.19}$ $85.13_{\pm 0.98}$ $85.02_{\pm 0.42}$
PhysioNet $74.98_{\pm 1.84}$ $74.29_{\pm 1.50}$ $75.07_{\pm 1.05}$ $\mathbf{76.01_{\pm 0.73}}$ $75.29_{\pm 1.21}$
SHHS $78.95_{\pm 0.92}$ $78.04_{\pm 1.21}$ $80.30_{\pm 1.29}$ $78.84_{\pm 0.43}$ $\mathbf{80.32_{\pm 0.91}}$
SOF $86.30_{\pm 0.40}$ $85.82_{\pm 0.22}$ $86.57_{\pm 0.60}$ $86.31_{\pm 0.27}$ $\mathbf{86.99_{\pm 0.33}}$
Mean(Dataset) $80.50_{\pm 0.51}$ $79.41_{\pm 0.73}$ $80.95_{\pm 0.36}$ $\mathbf{81.22_{\pm 0.20}}$ $81.19_{\pm 0.11}$
Mean(Subject) $80.05_{\pm 0.78}$ $79.09_{\pm 0.90}$ $81.31_{\pm 0.83}$ $80.59_{\pm 0.19}$ $\mathbf{81.39_{\pm 0.69}}$
Balanced@400 ABC $\mathbf{79.80_{\pm 0.34}}$ $77.86_{\pm 0.80}$ $78.36_{\pm 1.20}$ $79.49_{\pm 0.68}$ $78.08_{\pm 0.78}$
[…]
CFS $87.01_{\pm 0.18}$ $85.61_{\pm 0.16}$ $\mathbf{87.62_{\pm 0.27}}$ $87.37_{\pm 0.44}$ $87.06_{\pm 0.77}$
CHAT $66.56_{\pm 1.42}$ $61.32_{\pm 2.25}$ $64.19_{\pm 4.63}$ $69.90_{\pm 2.74}$ $\mathbf{71.86_{\pm 0.95}}$
[…]
MROS $83.69_{\pm 0.39}$ $82.22_{\pm 1.27}$ $83.95_{\pm 0.53}$ $\mathbf{84.15_{\pm 0.46}}$ $83.51_{\pm 0.84}$
PhysioNet $\mathbf{76.26_{\pm 1.27}}$ $70.40_{\pm 0.14}$ $73.84_{\pm 0.93}$ $75.24_{\pm 2.72}$ $73.51_{\pm 3.05}$
SHHS $76.98_{\pm 0.70}$ $75.98_{\pm 0.22}$ $79.12_{\pm 0.96}$ $78.19_{\pm 0.90}$ $\mathbf{79.26_{\pm 1.35}}$
SOF $85.49_{\pm 0.58}$ $84.23_{\pm 1.30}$ $85.50_{\pm 0.86}$ $\mathbf{85.56_{\pm 0.90}}$ $84.14_{\pm 1.05}$
Mean(Dataset) $9.64_{\pm 0.41} 77$$7.57_{\pm 0.73}$ $79.59_{\pm 0.25}$ $\mathbf{80.15_{\pm 0.26}}$ $80.12_{\pm 0.57}$
Mean(Subject) $78.57_{\pm 0.55}$ $76.86_{\pm 0.22}$ $79.53_{\pm 0.30}$ $79.70_{\pm 0.66}$ $\mathbf{80.29_{\pm 0.68}}$
A.8 Impact of Whitening and Target Covariance
@@
Dataset BatchNorm InstanceNorm PSDNorm
Barycenter Whitening
ABC $78.26_{\pm 1.33}$ $\mathbf{78.73_{\pm 0.42}}$ $8.18_{\pm 0.68} 77$$7.86_{\pm 1.33}$
CCSHS $87.42_{\pm 0.16}$ $87.62_{\pm 0.42}$ $87.58_{\pm 0.30}$ $\mathbf{87.80_{\pm 0.23}}$
CFS $84.32_{\pm 0.57}$ $\mathbf{84.72_{\pm 0.33}}$ $84.29_{\pm 0.36}$ $84.01_{\pm 0.60}$
CHAT $66.55_{\pm 0.88}$ $64.43_{\pm 4.41}$ $\mathbf{70.28_{\pm 1.70}}$ $69.07_{\pm 3.73}$
HOMEPAP $75.25_{\pm 0.50}$ $76.47_{\pm 0.63}$ $\mathbf{76.83_{\pm 0.61}}$ $76.13_{\pm 0.93}$
MASS $70.00_{\pm 1.91}$ $71.52_{\pm 1.13}$ $\mathbf{72.77_{\pm 1.09}}$ $69.11_{\pm 1.51}$
MROS $80.37_{\pm 0.20}$ $80.28_{\pm 0.21}$ $80.26_{\pm 0.11}$ $\mathbf{80.50_{\pm 0.75}}$
PhysioNet $\mathbf{75.81_{\pm 0.13}}$ $74.68_{\pm 0.55}$ $74.82_{\pm 2.11}$ $74.58_{\pm 1.57}$
SHHS $76.44_{\pm 0.92}$ $78.68_{\pm 0.37}$ $\mathbf{78.88_{\pm 0.68}}$ $78.77_{\pm 0.67}$
SOF $81.08_{\pm 1.14}$ $80.68_{\pm 1.38}$ $79.49_{\pm 0.41}$ $80.10_{\pm 0.62}$
Mean $7.55_{\pm 0.34} 77$$7.78_{\pm 0.46}$ $\mathbf{78.34_{\pm 0.42}}$ $77.79_{\pm 0.30}$
A.9 Generalization of PSDNorm in CNNTransformer